In a year, Seema earns ₹ 1,50,000 and saves ₹ 50,000. Find the ratio of
Money that she saves to the money she spends.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Money earned by Seema=₹150000
Money saved by her =₹50000
Money spent by her =₹150000 -₹50000 =₹100000
Ratio of money saved to money spent = 50000 / 100000 = 5 / 10
= 1/2

Thanks